identify the 4 major peaks and functional groups for purified caffeine on this spectra

1% 4000 3500 3000 2500 2000 1500 1000 450 cm-1
Answer #1

Alchol-OH- streching -C-H stretch 1% =C-H stretch CEO ketone C=0 amide 3500 3000 2500 2000 1500 1000 450 cm-1